# Limits to Predicting Online Speech Using Large Language Models

### v1

@@ A.1.3 Preprocessing tweets
Before our experiments, we preprocessed our collected tweets by filtering out urls, deduplicated spaces and fixed some special character encodings. We found that urls were not relevant in analysing the predictability of (organic) online speech, while deduplication of spaces is a fairly common preprocessing step in NLP. For our experiments described in Section 4.3, we further filtered out @-mentions and hashtags.
Appendix B Prompting
B.1 Individual variability
@@
B.2 Correlation with average tweet length
@@
(a) Llama-2
@@
B.3 Hard to predict users gain more from additional context
@@
B.4 Two control groups: social and temporal control
@@
B.5 “Surprise effect” of first token
LLMs are known to have high negative log-likelihood on the first token of the input sequence. This could artificially inflate model uncertainty in the no context setting. To test its effect on our results, we omit predictions on the first token from our analysis. Figure 20 illustrates the results of this experiment. Indeed, the gap between the no context and other context settings shrink as expected. However the differences still remain significant and thus don’t change our main findings.
Figure 20: The effect of taking the first token out of the analysis. Overall, the average BPC decreases (predictability goes up), however the most significant drop is in the no context settings for GPT-2-XL and Llama-2. This happens because frequently the NLL of the first token (with no token preceding it) is usually quite high. Lighter bars are results from our original experiment.
B.6 Most improved token due to random context: @-mention
We were interested in which tokens benefited most after including random (i.e. social) context. We selected tokens with >100 occurences, and ranked them based on how much their prediction improved on average. Table 1 shows that the ’@’ token benefits the most across all three models, indicating that @-mentions are locations of greatest improvement. See Table 2 for a full table on top-10 most improved token predictions (with all possible contexts).
@@ B.6 Most improved token due to random context: @-mention
Appendix C Finetuning
C.1 Pre-finetuning
(a) Loss calculated on the 5% validation split ($\sim 0.5$M tweets).
Figure 21: Loss curves of the pre-finetuning of GPT-2-XL on 10M tweets for 1 epoch. Validation loss (negative log-likelihood) was calculated on a 5% split. To make sure that pre-finetuning also improved subject predictability, we also selected a random subset ($n=100$) of the 5k subjects. We used their evaluation tweets to calculate the loss and since each $|\mathcal{T}_{u}^{\texttt{eval}}|=250$, this means a total of 25000 tweets.
GPT-2-XL Falcon Llama-2
social temporal peer user social temporal peer user social temporal peer user
0 @ @ hetti ĠKraft @ @ hyde 749  @  @ Extra  Kraft
1 âĢ âĢ DOM hetti Brit Brit perm Ġgenealogy  Hey  Ain DOM medium
2 Ļ Ļ 204 ĠARTICLE ĠAmen ĠNah Agg Ich  Ain  Hey member Extra
3 ĠARTICLE ĠARTICLE Agg hyde ĠNah ĠBru 018 ildo  Happy  tf zent zent
4 ĠðŁij ĠðŁij rium DOM ĠBru ĠAmen 641 hyde  Wait Fil $\upsilon$ $\upsilon$
5 ĠðŁ ĠâĢ alys ĠSag ĠSame ĠDamn atts MCA  tf  Wait Los  Via
6 ĠâĢ ĠðŁ perm medium ĠDamn ĠWait 576 perm /@ soft members hour
7 Ġâĺ ĠPis Hour ĠRescue ĠWait ĠSame 931 ĠKeller rach rach  @ member
8 Ġâľ ľ hyde ĠPis ĠDang ĠDude Chi ENV  Okay  Via gat DOM
9 ľ Ġâĺ Extra ĠKeller ĠNope ellan 454 ĠVenue  Ton  Om  txt ihe

### v2

@@ A.1.3 Preprocessing tweets
Before our experiments, we preprocessed our collected tweets by filtering out urls, deduplicated spaces and fixed some special character encodings. We found that urls were not relevant in analysing the predictability of (organic) online speech, while deduplication of spaces is a fairly common preprocessing step in NLP. These preprocessing choices were partially inspired by an open-source project called HuggingTweets Dayama (2022). For our experiments described in Section 4.2, we further filtered out @-mentions and hashtags.
(a) Loss calculated on the 5% validation split ($\sim 0.5$M tweets).
Figure 14: Loss curves of tweet-tuning GPT-2-XL on 10M tweets for 1 epoch. Validation loss (negative log-likelihood) was calculated on a 5% split. To make sure that tweet-tuning also improved subject predictability, we also selected a random subset ($n=100$) of the 5k subjects. We used their evaluation tweets to calculate the loss and since each $|\mathcal{T}_{u}^{\texttt{eval}}|=250$, this means a total of 25000 tweets.
A.2 Models
A.2.1 Training data
GPT-2 was trained on WebText with content up to December 2017 Radford et al. (2019) ($95\%$ of our subjects’ tweets were authored after 2022), Falcon was trained on RefinedWeb which is is built using all CommonCrawl dumps until the 2023-06 one Penedo et al. (2023) (CommonCrawl typically does not contain snapshots of Twitter com (2024)) and the authors of the Llama models claim to have “made an effort to remove data from certain sites known to contain a high volume of personal information about private individuals.” Touvron et al. (2023); Dubey et al. (2024). Llama-2 has a knowledge cutoff at September 2022, while Llama-3 8B has seen data up to March 2023. Based on these facts we believe there is limited risk of data contamination.
A.2.2 Tweet-tuning on 10M tweets
We finetuned some models on the 10M tweets we collected during the sampling stage of our data collection process (as described Section 3.1), where a 5% split was reserved for validation. Tweets were concatenated, with the special eos token ’<|endoftext|>’ serving as a separator between them.
GPT-2-XL-tt
We finetuned all parameters for 1 epoch, using a constant learning rate of $5\mathrm{e}{-5}$, batch size 8 and fp16 mixed precision training on a single A-100 80GB GPU with an AdamW optimizer. We used the example finetuning script from the transformers library as our base222run_clm.py from here: https://github.com/huggingface/transformers/tree/main/examples/pytorch/language-modeling, where we kept most of the default training arguments.
Llama-3-8b-tt
We finetuned all parameters for 5 epochs, using an initial learning rate of $2\mathrm{e}{-5}$ and a cosine learning rate scheduler. We used a paged AdamW optimizer with batch size 1 and gradient accumulation steps set to 8. We used axolotl for finetuning, the base for our config was their example full finetune script [^3] https://github.com/axolotl-ai-cloud/axolotl/blob/main/examples/llama-3/fft-8b.yaml for Llama-3-8b (only the initial learning rate and number of epochs was changed).
A.3 Prompting experiments
Let us denote a single tweet as $T=(t_{1},t_{2},...t_{m})$, where $t_{i}$ is a single token ($i=0...m$). A token is an item in the LLM’s vocabulary, and can be thought of as a collection of characters that frequently co-occur. Each tweet has a maximum of 280 characters [^4] This limit changed to 4000 characters on the 09. February 2023. and after tokenization most tweets have between 0-100 tokens (Fig. 20). We use a language model with parameters $\theta$ to predict token $t_{i}$ based on the preceding tokens $t_{<i}$. The model’s output will be a likelihood over all possible tokens in the model’s vocabulary, however we are only interested in the conditional probability of $t_{i}$: $p_{\theta}(t_{i}|t_{<i})$.
Calculating negative log-likelihood
$\bar{L}_{u}=\frac{1}{n}\sum_{T_{j}\in\mathcal{T}^{\texttt{eval}}_{u}}L(T_{j}),$
where $n$ is the total number of tokens. We additionally introduce notation to distinguish what context was used to calculate: $\bar{L}_{u}^{c}$, where subscript $c\in\{\texttt{user},\texttt{peer},\texttt{random}\}$. Here, the conditional probability of token $t_{i}$ is based on preceding tokens $t_{<i}$ as well as tokens from the appropriate context: $p_{\theta}(t_{i}|\mathcal{T}_{u}^{c},t_{<i})$. The added context lends a convenient cross-entropy like interpretation of the shared information between the context and evaluation tweets we are trying to predict.
Sensitivity to prompting strategy
Tweets that served as context were concatenated using the ‘newline’ token for GPT-2-XL and the ‘space’ token for the other models (they had no standalone ‘newline’ token like GPT-2). Using low-frequency tokens as a separator between tweets (such as the eos token, which is usually reserved for separating training documents) produced abnormally high NLLs, which is why we decided to use more common tokens, such as space or newlines. We used the following input sequence lengths: 1024 tokens for GPT-2-XL, 2048 for Falcon 40B and 4096 for Llama-2 and Llama-3 to calculate the token probabilities. In case the provided context exceeded this length, the oldest tweets were discarded.
Conversion to bits per character
$\overline{bpc}_{u}=\bar{L}_{u}\cdot\frac{1}{\bar{C}_{u}}\cdot\frac{1}{\ln 2}\,.$
Estimating the effect size of context on predictability
A.4 Finetuning experiments
Tweet-tuned models were additionally finetuned on one of the contexts. We also experimented with using mixtures of different contexts. Here, our proxy for unpredictability was the negative log-likelihood (or also commonly known as the cross-entropy loss) in the final round of fine-tuning. Tweets were concatenated, with the special eos token ’<|endoftext|>’ serving as a separator between them. Reported results are calculated on $\mathcal{T}^{\texttt{eval}}_{u}$.
We only ran this experiment on the tweet-tuned version of GPT-2 (GPT-2-XL-tt), because of the non-negligible amount of time and resources this experiment requires to run. Per subject there are 3 different types of context, plus 4 mixtures of contexts. That is, we finetuned GPT-2 a total of $7*5102=35714$ times. Together with tweet-tuning, it cost us approximately 1.6 e19 FLOPs to run this experiment.  [^6] As a comparison: it took OpenAI around e20-e21 FLOPs to finetune GPT-2 from scratch. While finetuning on larger models would certainly result in lower cross-entropy overall, we do not believe that it would change our main conclusions. This is supported by the high agreement on the ranking of users across models (Fig. 16) as well as the stability of our conclusions for our prompting experiments as we go up in model size.
Appendix B Supplementary Results
B.1 High agreement on user ranking
B.2 Individual variability
@@
B.3 Correlation with average tweet length
@@
Figure 19: Average improvement in NLL from additional user context (compared to none). The deterioration of predicting the first token correctly is an artifact of how we ran our experiments. Model: Llama-3-8B.
(a) GPT-2-XL
@@
B.4 Hard to predict users gain slightly more from additional context
@@
B.5 Two control groups: social and temporal control
@@
B.6 First token uncertainty
Adding context can help in correctly predicting the first token. This is the case for GPT-2, Falcon and Llama-2 models (see Fig. 5), where models learn from context that a tweet often begins with an @-mention. For Llama-3, we see a slightly different picture in Figure 19: predicting the first token correctly becomes harder because of the added context. Upon investigating this difference, we found that this is simply an artifact of how we ran our experiments. Llama-3 assigns much higher probability to predicting ’@’ as the first token to begin with. With the added context however, it assigns high probability to ’ @’ (a space followed by an @), which is encoded as a single token in Llama-3 ([571]). Because of how we ran our experiments, the space ([31]) and the subsequent @-mention ([220]) got encoded separately, resulting in tokens [31, 220] instead of the expected [571]. Thus we decided to exclude the first token from our analysis when reporting results on Llama-3.
Figure 23 illustrates how dropping the first token from our analysis affects model uncertainty across all four models. Unsurprisingly, model uncertainty goes down. For models which had high negative log likelihoods on the first token improved the most in the no context setting. On the other hand – for reasons we outlined above – model uncertainty on the Llama-3 model improved the most outside the no context setting. Overall, by excluding the first token from our analysis, we found that our main results from Figure 2 are nicely replicated on Llama-3 as well.
Figure 23: The effect of taking the first token out of the analysis. Overall, the average model uncertainty decreases (predictability goes up). Lighter bars are results from our original experiment.
B.7 Most improved token due to random context: @-mention
We were interested in which tokens benefited most after including random (i.e. social) context. We selected tokens with >100 occurences, and ranked them based on how much their prediction improved on average. Table 1 shows that the ’@’ token benefits the most across all three models, indicating that @-mentions are locations of greatest improvement. See Table LABEL:tab:top_10_full for a full table on top-10 most improved token predictions (with all possible contexts).
@@ B.7 Most improved token due to random context: @-mention
B.8 Falcon-40b w/o @-mentions and hashtags
We repeated our prompting experiments after removing @-mentions and hashtags from our tweets (see Figure 6). An interesting observation for our results on Falcon is that after removing @-mentions and hashtags, additional context did not improve predictability the same way as it did before. In case of peer and random context, it even increased model uncertainty. The effect size of the user context dropped to only $0.4\sigma$, which is significantly less than the $2.1\sigma$ from before. A possible explanation is that the only (useful) predictive signal Falcon was able to pick up on was in the removed pieces of text. Another possibility is that it is simply more sensitive to discontinuities inside the text than other models.
Figure 24: Model uncertainty on a random subject. Increasing the context window size lowers model uncertainty. Dashed lines are model uncertainties reached with 4096 context window size, with the specified context. Model: Llama-2-70b.
B.9 Negative log-likelihood convergence
We illustrate how negative log-likelihood converges as we increase the context window size. In other words, we show that the more tokens (tweets) we include, the lower NLL we get. We concatenated tweets in $\mathcal{T}_{u}^{\texttt{eval}}$ for a random user $u$ using the ’space’ token as a separator. In Figure 24, we show how NLL converges as we increase the context window size from 10 tokens to 4096 tokens. In the end, it converges to the same NLL we reach when we condition on concatenated tweets from the user context $\mathcal{T}_{u}^{\texttt{user}}$. As we can see, LLMs may need a fairly long context size for the negative log-likelihood to converge.
GPT-2-XL
social temporal peer user
1 @ @ hetti ĠKraft
2 âĢ âĢ DOM hetti
3 Ļ Ļ 204 ĠARTICLE
4 ĠARTICLE ĠARTICLE Agg hyde
5 ĠðŁij ĠðŁij rium DOM
6 ĠðŁ ĠâĢ alys ĠSag
7 ĠâĢ ĠðŁ perm medium
8 Ġâĺ ĠPis Hour ĠRescue
9 Ġâľ ľ hyde ĠPis
10 ľ Ġâĺ Extra ĠKeller
Falcon
social temporal peer user
1 @ @ hyde 749
2 Brit Brit perm Ġgenealogy
3 ĠAmen ĠNah Agg Ich
4 ĠNah ĠBru 018 ildo
5 ĠBru ĠAmen 641 hyde
6 ĠSame ĠDamn atts MCA
7 ĠDamn ĠWait 576 perm
8 ĠWait ĠSame 931 ĠKeller
9 ĠDang ĠDude Chi ENV
10 ĠNope ellan 454 ĠVenue
Llama-2
social temporal peer user
1  @  @ Extra  Kraft
2  Hey  Ain DOM medium
3  Ain  Hey member Extra
4  Happy  tf zent zent
5  Wait Fil $\upsilon$ $\upsilon$
6  tf  Wait Los  Via
7 /@ soft members hour
8 rach rach  @ member
9  Okay  Via gat DOM
10  Ton  Om  txt ihe